OpenStudy (anonymous):

integrate [csch^2(2/x)]/(x^2) dx I used 'u' substitution with x^-2 which gives me -1/2du=dx/x^3 I know the answer is 1/2coth(2/x)+C so I think I'm pretty close, I just can't find the 'du' in the integral with the substitution I have....where did I go wrong?

hartnn (hartnn):

put u= 2/x du=.... ?

OpenStudy (anonymous):

Without writing that out, I don't think I'll end up with a du to sub back in to put in terms of u..

hartnn (hartnn):

means ? if u=2/x du = -2/x^2 dx so, (1/x^2) dx = du/(-2) substitute this in your original integral.

OpenStudy (anonymous):

wouldn't it be (-1/2)du=(1/x^2)dx?

OpenStudy (anonymous):

Didn't see your / ... nevermind my stupidity. Thanks

hartnn (hartnn):

ok..so now can you solve ?

OpenStudy (anonymous):

Yeah, got it now...didn't see (2/x) working in my head...I should really learn to write/test things out. Appreciate the help!

hartnn (hartnn):

welcome ^_^